# ILRe: Intermediate Layer Retrieval for Context Compression in Causal Language Models

### v1

@@ 4.4 Efficiency
In Figure 1, we measure the Time To First Token (TTFT) of our approach with a batch size of one on a single Huawei Ascend 910B NPU, averaging results across multiple trials. Due to hardware limitations, we only measured the full-context baseline up to $256K$ tokens, and extended the results to $1M$ tokens using a quadratic fit. Given the large discrepancy in TTFT between our approach and the full-context baseline, we use the log-scale y-axis in the figure. Our method demonstrates a significant speedup, achieving up to a hundredfold improvement in long-context inference. Additionally, the linear fit of ILRe yields an $R^{2}$ value of $0.994$, which strongly aligns with our complexity analysis conclusion of $O(L)$ in Section 3.4.
@@ 5 Conclusion
In this study, we propose a training-free context compression framework, ILRe. Built upon standard causal LLMs, it is highly efficient and effective in context prefilling and extraction. By integrating streaming chunked prefill, an observation window, multi pooling kernels and an early exiting layer, ILRe reduces the computational overhead of self-attention, delivering up to a hundredfold speedup in end-to-end inference while maintaining remarkable accuracy. Our experimental evaluations demonstrate that ILRe holds significant promise for advancing long-context LLM inference.

### v2

@@ 4.4 Efficiency
In Figure 1, we measure the Time To First Token (TTFT) of our approach with a batch size of one on a single Huawei Ascend 910B NPU, averaging results across multiple trials. Due to hardware limitations, we only measure the FullContext baseline up to $256K$ tokens, and extend the results to $1M$ tokens using a quadratic fit. Given the large discrepancy in TTFT between our approach and the full-context baseline, we use the log-scale y-axis in the figure. Our method demonstrates a significant speedup, achieving up to a hundredfold improvement in long-context inference. Additionally, the linear fit of ILRe yields an $R^{2}$ value of $0.994$, which strongly aligns with our complexity analysis conclusion of $O(L)$ in Section 3.4.
@@ 5 Conclusion
In this study, we propose a training-free context compression framework, ILRe. Built upon decoder-only causal LLMs, it is highly efficient and effective in context prefilling and extraction. By integrating streaming chunked prefill, an observation window, multi pooling kernels and an early exiting layer, ILRe reduces the computational overhead of self-attention, delivering up to a hundredfold speedup in end-to-end inference while maintaining remarkable accuracy. Our experimental evaluations demonstrate that ILRe holds significant promise for advancing long-context LLM inference and length extrapolation.
@@ A.1 Details of Retrieval Task
Similar to Figure 2, the results of Llama-3.1-8B-Instruct and Qwen2.5-7B-Instruct in the Retrieval Task are present in Figure 4 and  5 respectively.
(a) Different layers.
(a) Different depths.
Figure 4: Retrieval Task results of model Llama-3.1-8B-Instruct over $128K$ context.
(a) Different layers.
(a) Different depths.
Figure 5: Retrieval Task results of model Qwen2.5-7B-Instruct over $32K$ context.
A.2 Occurrence of Max Attention Heads
In Figure 6, the statistics of the maximum attention head indices in the retrieval task are present. This data is aggregated after the max operation of attention matrix over the query dimension.
Figure 6: Occurrence Ratios of max attention head indices in one $128K$ Retrieval Task of model Llama-3.1-Ultralong-8B-1M-Instruct.
A.3 Setups of benchmarks
[…]
For LongBench, we utilize the toolset introduced in (Yuan et al., 2024), and extract the last 64 tokens from each input prompt as the query (also with the remaining tokens serving as context). Given that LongBench contexts are relatively short (with an average length of fewer than $16K$ tokens), we only test it with two short length capacity models: Llama-3.1-8B-Instruct (Grattafiori et al., 2024) and Qwen2.5-7B-Instruct (Qwen et al., 2025).